# State Frequency Estimation for Anomaly Detection

### v1

@@ Abstract
Many works have studied the efficacy of state machines for detecting anomalies within NetFlows. These works typically learn a model from unlabeled data and compute anomaly scores for arbitrary traces based on their likelihood of occurrence or how well they fit within the model. However, these methods do not dynamically adapt their scores based on the traces seen at test time. This becomes a problem when an adversary produces seemingly common traces in their attack, causing the model to miss the detection by assigning low anomaly scores. We propose SEQUENT, a new approach that uses the state visit frequency to adapt its scoring for anomaly detection dynamically. SEQUENT subsequently uses the scores to generate root causes for anomalies. These allow the grouping of alarms and simplify the analysis of anomalies. Our evaluation of SEQUENT on three NetFlow datasets indicates that our approach outperforms existing methods, demonstrating its effectiveness in detecting anomalies.
I Introduction
NetFlow encapsulates summarized network statistics of a connection between two hosts and can be used for traffic analysis. Studies have shown the effectiveness of state machines in detecting anomalies on NetFlows in an unsupervised manner [24, 10, 12]. Typically, these studies learn a model from benign flows and then compute an anomaly score based on the likelihood of a trace given the model.
@@ I Introduction
Although a standard approach in machine learning (ML), we argue that this has two limitations making it not well-suited for detecting anomalies we see in practice. Firstly, many benign flows occur with a low probability, resulting in a large number of false alarms. Secondly, malicious flows tend to resemble common flows, giving them large probabilities and thus making it hard to raise alarms. In this paper, we propose a way to overcome these limitations by using a key feature of the state machine model: its behavioral structure. Instead of computing likelihoods, we compute visit frequencies of different states within the model. Each state essentially represents a distinct type of network traffic behavior (a cluster or latent value). When we observe behavior much more frequently than expected, we raise an alarm. Intuitively, benign flows with small probability thus do not raise alarms when they occur as frequently as expected. Moreover, malicious flows with high probability do raise alarms when they occur more frequently than expected. Note that this also detects infrequent malicious flows that were not observed at training since their expected count is assumed to be very small. The price to pay is that now normal flows raise alarms when they occur more frequently. Although this does happen in network traffic, we believe and demonstrate in experiments that this is a price worth paying.
@@
We introduce SEQUENT, an approach that learns a state machine from benign NetFlows to model sequential network behavior. The process begins by discretizing the NetFlow features into a symbolic representation. We develop a discretization method [3] inspired by the Glove algorithm [25]. We discretize each selected feature independently and then map each NetFlow to a symbol, which is the concatenation of the individual discretized symbols.
@@
SEQUENT then runs each trace observed at test time on the learned model and computes the set of states visited by the trace together with their corresponding visit frequency. SEQUENT compares the visit frequencies to the expected frequencies observed at train time and raises an alarm whenever a type of behavior occurs more frequently than expected.
@@
A well-known challenge faced by many security analysts in their daily routine is the large volume of alerts raised by the anomaly detection system (ADS) [21, 14, 1], turning the inspection of alerts to a labor-intensive task. The sheer volume of alerts makes it difficult for analysts to prioritize which anomalies to inspect first. SEQUENT addresses this issue by computing root causes (state identifiers) for each detected anomaly, which allows grouping and ranking of anomalies (e.g., using the frequency of a root cause). Additionally, SEQUENT establishes links between the root causes and their corresponding NetFlows. The aim is to provide analysts with concrete instances of anomalous behavior, enabling a quick and in-depth analysis of anomalies.
@@
We empirically demonstrate SEQUENT’s effectiveness in the task of detecting various types of network anomalies using three publicly available NetFlow datasets [9, 11, 17]. Our experiment results demonstrate that SEQUENT is highly effective in the detection of network anomalies, outperforming existing state-of-the-art unsupervised methods. The main contributions of this paper are summarized as:
@@
II Preliminaries
II-A State Machine
@@
A state machine, formally known as a finite state machine or finite state automaton, is a mathematical model used for modeling the sequential behavior of a system. It is formally defined as a 5-tuple $(\Sigma,Q,q_{0},\delta,F)$[26], where:
  • $\Sigma$ is a finite alphabet
  • $Q$ is a finite set of states
  • $q_{0}\in Q$ is a unique start state
  • $\delta:Q\times\Sigma\to Q$ is the transition function
  • $F\subseteq Q$ is the set of final states
@@
Given an input trace $t$ (a sequence of events, where each event $e\in\Sigma$) and a system $S$, one can use the state machine to determine in which state $S$ would end up, thereby understanding the behavior the system would exhibit. An illustrative example of a state machine is shown in Figure 1. The state machine in this example consists of only one final state, $q_{3}$.
@@
Figure 1: An illustrative example of a state machine. At the end of the input trace, $aabab$, we would end up in state $q_{2}$.
II-B State Machines Considered in SEQUENT
State machines come in two flavors, namely deterministic or non-deterministic state machines [26]. SEQUENT solely considers deterministic state machines as they are easier to interpret; simpler to predict the next event, and verifying the same sequence of events would execute the same sequential behavior is more straightforward [15]. We can employ two approaches for learning a state machine from trace data: active and passive learning. We opt for the latter approach as it only requires us to collect traces from the target system $S$ instead of continuously interacting with $S$ to learn a state machine [2].
@@ II-B State Machines Considered in SEQUENT
State-merging heuristic algorithm is a highly effective method for learning state machines in the passive-learning approach [13]. It first starts with a tree-like state machine (see Figure 2) known as the Prefix Tree Acceptor (PTA). It then iteratively compares pairs of states to identify states that could be merged. Merges are done based on the scores computed during the comparisons. The algorithm ends when it cannot find any new merges. The final output is an approximation of the smallest state machine consistent with the input data.
Figure 2: An example of a PTA.
III Assumptions of the Adversary
@@
Our work is grounded in a set of assumptions we formulated concerning the objectives and capabilities of an adversary. We consider an adversary attacking an arbitrary target network $TN$, driven by various objectives such as infiltrating $TN$ to find machines containing sensitive data, disrupting normal operations by launching denial of service attacks against specific services, orchestrating a botnet, or infecting machines with malware. Our main experiments take the train and test data from the datasets to simulate an adversary attacking $TN$ without taking our defense into account. In our robustness experiments, we assume a more adept adversary that is aware of an ADS within $TN$ but unaware of its specifics. Their goal is to remain undetected until they have accomplished their objectives. To accomplish this, the adversary can monitor $TN$ to collect normal network behaviors and then attempt to bypass the ADS by generating behaviors resembling the ones that were collected. The adversary can only listen to the data that was sent but cannot delete any of the sent data.
IV Related Works
@@
Traditional Methods. Several works have studied the efficacy of unsupervised learning for network anomaly detection. The primary focus has been on utilizing traditional unsupervised ML algorithms (e.g., Isolation Forest). Zoppi et al. conducted multiple studies evaluating the effectiveness of traditional unsupervised methods for anomaly detection on NetFlow data [31, 32]. In contrast to SEQUENT, these studies do not dynamically adapt the anomaly scores based on the events/traces seen at test time for anomaly detection.
@@
State Machines. Numerous works have investigated the efficacy of state machines for anomaly detection. Pellegrino et al. presented BASTA, a system that learns probabilistic timed state machines from NetFlows and uses them as behavioral fingerprints for machines infected with botnet malware [24]. Alerts are raised when traces match behaviors captured in the fingerprints. Unlike SEQUENT, BASTA uses malicious data for learning as opposed to solely benign data.
@@
Lin et al., Cao et al., and Matoušek et al. learned probabilistic state machines to detect anomalies in different systems[10, 16, 18]. These works raise alerts when a new trace deviates significantly from what was observed during training. The degree of deviation is quantified based on the likelihood of a trace. In contrast, SEQUENT quantifies the abnormality based on state visit frequency for a corresponding trace and compares this to what was observed during training.
@@
Grov et al. presented a different approach for learning state machines to detect anomalies in NetFlows [12]. The approach learns state machines using solely traces of network protocols extracted from benign data. The degree of abnormality is based on the normalized edit distance of an arbitrary trace under scrutiny; the number of changes needed such that each trace event can be replayed in the state machine. In contrast to SEQUENT, this work does not utilize the frequency of the states to adapt the scores dynamically at test time. Furthermore, it solely considers network protocols for the learning, whereas SEQUENT considers other features.
@@
Neural Networks. Deep Neural Networks (DNN) gained increasing popularity among researchers in recent years due to their ability to identify complex hidden patterns that traditional models may struggle to detect. Mirsky et al. introduced Kitsune, a state-of-the-art solution that utilizes an ensemble of Autoencoders (AE) to detect network anomalies in an unsupervised manner. It extracts various statistics from benign data and uses them for training. Kitsune then attempts to recreate instances it has seen during training by computing similar statistics for arbitrary new data. Alarms are raised for instances with large recreation errors[19]. Neural networks are notoriously known to be difficult to interpret, as it is not always clear what happens between the input and output layers. Nguyen et al. sought to address this issue with their framework, GEE. It uses a Variational Autoencoder (VAE) to detect anomalies and generates explanations by providing SHAP-like plots of the gradients computed for NetFlow features[23]. However, it is not clear how a security analyst should interpret these gradients. In contrast, SEQUENT learns an inherently interpretable model and generates root causes, which can be used to facilitate the understanding of (malicious) sequential behaviors. Root causes are linked back to concrete instances producing the anomalous behaviors.
V SEQUENT
V-A Intuition Behind SEQUENT
@@
SEQUENT employs a novel approach for computing the anomaly score of an arbitrary trace under scrutiny. Unlike the measures of abnormality presented in prior works [24, 18, 16, 12, 10], SEQUENT leverages the frequencies of the state visits in a state machine to compute an adaptive anomaly score. The underlying principle of SEQUENT is that behaviors occurring significantly more often than those observed during training are indications of anomalies. Traces under scrutiny are replayed on a state machine to determine how well they fit within the learned model. Each trace visits a set of states, and the frequency of visits to each state is compared to the expected frequency observed during training. When the frequency of visits to a state is significantly higher than its expected frequency, it signals that unusual behavior is occurring in this state. The more significant the increase in visits for this state, the more likely it is to show anomalous behavior. SEQUENT employs this rationale to dynamically adapt its scores for arbitrary traces seen at test time. SEQUENT’s intuition will also work for (anomalous) behaviors never seen before during training, as these will visit unusual paths during the replay.
V-B SEQUENT’s Anomaly Detection Pipeline
@@
Figure 3 depicts SEQUENT’s anomaly detection pipeline. It offers a high-level overview showing the flow of processing NetFlows, learning of a state machine, and finally generating the results. Each step is described in the following paragraphs.
Figure 3: SEQUENT’s anomaly detection pipeline.
@@
Processing NetFlows. NetFlow needs to be discretized before learning a state machine, given that it contains both continuous and discrete feature data. Each flow needs to be transformed into an event to generate traces. The discretization process is not strictly necessary for the learning of a state machine, but using the raw NetFlow feature data (specifically continuous data) will lead to the combinatorial explosion of states. Discretization enables us to lower the number of possible combinations in the feature data, reducing the number of states used in the state machine. To discretize NetFlows, we develop an encoding algorithm inspired by the GloVe algorithm [25]. Our encoding algorithm computes co-occurrences between values for a given flow feature $f$ and stores them in a matrix. Each row denotes a unique value $v$ of $f$ and how often it occurs with another value $v^{\prime}$. Feature values are discretized by first clustering the rows using KMeans and then using cluster assignments as their discretization. We make our encoding algorithm available on Github [3]. To transform a flow into an event, we concatenate the discretized feature values with an underscore symbol to form a single event symbol. For example, if three features are used, we get the following event symbol: $DFV1\_DFV2\_DFV3$, where $DFV$ is the abbreviation for discretized feature value. In the case of one feature (e.g., bytes), we directly use its discretized values. To convert events into traces, we apply a sliding window of length $N$ to generate traces. These are then subsequently used to learn a state machine. Traces that do not have a length of $N$ are excluded from learning as they do not provide sufficient information.
@@
Learning of a State Machine. SEQUENT utilizes FlexFringe [28] to learn a state machine from NetFlow traces. This framework offers, out of the box, various heuristics for learning a state machine. Additionally, it enables its users to learn other types of models e.g., Markov Chains. Furthermore, FlexFringe automatically creates traces from CSV files provided as input data. The user has to specify which features to use to create events, the length of the sliding window, and how the data should be sorted (e.g., by connections) to create the traces. We opt for FlexFringe based on these aforementioned features.
Computing Rolling Anomaly Score. Following the intuition described in Subsection V-A, we construct a formula for computing a rolling anomaly score for a state $q_{i}\in Q$ visited by an arbitrary trace $t$:
anomaly score (qi)=observations of qi so $\textit{anomaly score $(q_{i})$}=\frac{\sum\textit{observations of $q_{i}$ so far}}{E}$
where $E$ is defined as:
$E=\frac{\sum\ \textit{observations of $q_{i}$ in training traces}}{\sum_{i=0}^{n}\ \textit{observations of $q_{i}$ in training data}}\cdot\textit{UC}$
Then, UC is defined as the unique number of states seen so far in new traces:
$UC=\sum_{j=0}^{m}|StateSequence(t_{m})|$
and $StateSequence$ is a function that returns a set of states visited by a trace $t$:
$StateSequence(t)\rightarrow\{(q_{1},q_{2},...,q_{n})\},\textit{where }q_{i}\in Q$
UC only concerns the set of states learned during training, i.e., no new states are learned at test time. As $t$ can visit a particular state $q_{i}$ multiple times (e.g., if $q_{i}$ is part of a loop within the state machine model), this could skew the observations of $q_{i}$ (thus also the abnormality of $q_{i}$) for subsequent traces. To reduce this impact, we take the logarithm of the anomaly score computed for $q_{i}$, resulting in the following formula:
anomaly score ($\textit{anomaly score $(q_{i})$}=\log\left(\frac{\sum\ \textit{observations of$q_{i}$ so far}}{E}\right)$
This formula enables us to change the level of abnormality for a state $q_{i}$ based on how frequently it is visited by each $t$ seen at test time. The anomaly score for $t$ is the sum of the anomaly scores computed for the set of states it has visited.

### v2

@@ Abstract.
Many works have studied the efficacy of state machines for detecting anomalies within NetFlows. These works typically learn a model from unlabeled data and compute anomaly scores for arbitrary traces based on their likelihood of occurrence or how well they fit within the model. However, these methods do not dynamically adapt their scores based on the traces seen at test time. This becomes a problem when an adversary produces seemingly common traces in their attack, causing the model to miss the detection by assigning low anomaly scores. We propose SEQUENT, a new unsupervised approach that uses the state visit frequency of a state machine to adapt its scoring dynamically for anomaly detection. SEQUENT subsequently uses the scores to generate root causes for anomalies. These allow the grouping of alarms and simplify the analysis of anomalies. We evaluate SEQUENT’s effectiveness in detecting network anomalies on three publicly available NetFlow datasets and compare its performance against various existing unsupervised anomaly detection methods. Our evaluation shows promising results for using the state visit frequency of a state machine to detect network anomalies.
NetFlow, Anomaly Detection, Unsupervised Learning, Automata Learning, Machine Learning
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; XXX - XXXX, 2025; XXXccs: Computing methodologies Anomaly detection
1. Introduction
NetFlow encapsulates summarized network statistics of a connection between two arbitrary hosts and can be used for traffic analysis. It is a popular alternative to storing individual packets collected from a network due to its benefits: easier to collect, simpler to process, and less intrusive to the privacy of the network users as there is no possibility to inspect every packet. Despite these benefits, detecting network anomalies (e.g., network attacks) within NetFlows (or simply flows) remains a challenge, as essential information might be lost within the summarized statistics. Various works have studied the use of supervised and unsupervised machine learning (ML) models to detect network anomalies within NetFlow data (Nguyen et al., 2019; Zoppi et al., 2021a; Venturi et al., 2021; Fosić et al., 2023; Hassan et al., 2024; Zoppi et al., 2021b; Campazas-Vega et al., 2023).
While there has been success in applying ML models to detect anomalies in NetFlow data, a challenge faced by many security analysts is the interpretability of the predictions made by the ML model (Nadeem et al., 2023). Understanding these predictions enables analysts to determine whether the model has made the correct decision and to gain confidence in the ML model deployed in the ADS.
To address this challenge, several studies have explored the use of state machines to detect network anomalies within NetFlow data, as these models are considered inherently interpretable (Pellegrino et al., 2017; Cao et al., 2022; Grov et al., 2019). Typically, these studies train the model using traces generated from benign flows and compute an anomaly score based on the likelihood of those traces according to the model.
@@ 1. Introduction
Although a standard approach in ML, we argue that this has two limitations, making it not well-suited for detecting anomalies we see in practice. Firstly, many benign flows occur with a low probability, resulting in many false alarms. Secondly, malicious flows tend to resemble common flows, giving them large probabilities and thus making it hard to raise alarms. To address this, we introduce SEQUENT (short for State Frequency Estimation for Anomaly Detection) to overcome these limitations by using a key feature of the state machine model: its behavioral structure. Instead of computing likelihoods, we compute visit frequencies of different states within the model. Each state essentially represents a distinct type of network traffic behavior (a cluster or latent value). When we observe behavior that occurs much more frequently than expected, an alarm is raised. Intuitively, benign flows with low probability do not trigger alarms if they occur at the expected frequency. Moreover, malicious flows with high probability do raise alarms when they occur more frequently than expected. Note that this also detects infrequent malicious flows not observed at training since their expected count is assumed to be very small. The drawback is that normal flows may raise alarms when they occur more frequently. While this can happen, our experiments demonstrate that these flows have a smaller impact compared to the malicious flows.
@@
Similar to previous studies that utilized state machines, SEQUENT learns a state machine from benign flows to model sequential network behavior. The process begins by discretizing the flow features into a symbolic representation. To achieve this, we developed a discretization method (Cao, 2022a) inspired by the GloVe algorithm (Pennington et al., 2014). Each selected feature is discretized independently and each flow is mapped to a symbol, created by concatenating the discretized feature values of the corresponding flow.
@@
SEQUENT then runs each trace observed at test time through the learned model and computes the set of states visited by the trace. It uses the computed sets to update the visit frequency of each state in the model. SEQUENT compares the visit frequencies to the expected frequencies observed during training and raises an alarm whenever a behavior occurs more frequently than expected.
@@
Another well-known challenge faced by many security analysts in their daily routine is the large volume of alerts raised by the anomaly detection system (ADS) (Nadeem et al., 2022; Lanvin et al., 2023; Alahmadi et al., 2022). The sheer volume of alerts makes it difficult for analysts to prioritize which anomalies to inspect first, turning the inspection of alerts into a labor-intensive task. SEQUENT addresses this issue by computing root causes (state identifiers) for each detected anomaly, which allows grouping and ranking of anomalies (e.g., using the frequency of a root cause). Additionally, SEQUENT establishes links between the root causes and their corresponding flows. The aim is to provide analysts with concrete instances of anomalous behavior, enabling a quick and in-depth analysis of anomalies.
@@
We empirically demonstrate SEQUENT’s effectiveness in the task of detecting various types of network anomalies using three publicly available NetFlow datasets (Cao and Blaise, 2022; Garcia et al., 2014; Maciá-Fernández et al., 2018). Our experiment results demonstrate that SEQUENT is highly effective in the detection of network anomalies, outperforming existing state-of-the-art unsupervised methods. The main contributions of this paper are summarized as:
@@
2. Preliminaries
2.1. State Machine
@@
A state machine, formally known as a finite state machine or finite state automaton, is a mathematical model used for modeling the sequential behavior of a system. It is formally defined as a 5-tuple $(\Sigma,Q,q_{0},\delta,F)$(Sipser, 2013), where:
  • $\Sigma$ is a finite alphabet.
  • $Q$ is a finite set of states.
  • $q_{0}\in Q$ is a unique start state.
  • $\delta:Q\times\Sigma\to Q$ is the transition function that determines which transition should be taken based on the current state and the (new) symbol read from the input.
  • $F\subseteq Q$ is the set of final states.
@@
Given an input trace $t$ (a sequence of events, where each event $e\in\Sigma$) and a system $S$, one can use the state machine to determine in which state $S$ would end up, thereby understanding the behavior the system would exhibit. In the context of this work, each event in a trace represents a single flow. An illustrative example of a state machine is shown in Figure 1. The state machine in this example consists of only one final state, $q_{3}$.
@@
Figure 1. An illustrative example of a state machine. At the end of the input trace, $aabab$, we would end up in state $q_{2}$.
2.2. State Machines Considered in SEQUENT
State machines come in two flavors, namely deterministic or non-deterministic state machines (Sipser, 2013). SEQUENT solely considers deterministic state machines as they are easier to interpret; simpler to predict the next event, and verifying the same sequence of events would execute the same sequential behavior is more straightforward (Lee, 2021). With a non-deterministic state machine, multiple transitions can be triggered simultaneously, even when no event has occurred. These non-deterministic transitions can inaccurately increase the visit frequencies of states within a state machine, causing states (or behaviors) to be falsely flagged as anomalous. Since SEQUENT aims to address the challenge of large volume alerts by providing root causes for anomalies, non-deterministic transitions can result in providing false root causes to analysts.
Furthermore, the use of state machine is intended provide analysts with a way to inspect the sequential behaviors exhibited by an arbitrary system. However, when multiple transitions are triggered simultaneously within a non-deterministic state machine, it becomes difficult for an analyst to determine which path should be followed in the model given an arbitrary trace. For these reasons, SEQUENT focuses only on learning deterministic state machines.
@@ 2.2. State Machines Considered in SEQUENT
We can employ two approaches for learning a state machine from trace data: active and passive learning. We opt for the latter approach as it only requires us to collect traces from the target system $S$ instead of continuously interacting with $S$ to learn a state machine (Angluin, 1987). State-merging heuristic algorithm is a highly effective method for learning state machines in the passive-learning approach (Lang et al., 1998). It begins with a tree-like state machine, known as the Prefix Tree Acceptor (PTA) (see Figure 2). It then iteratively compares pairs of states to identify states that can be merged. Merges are done based on the scores computed during the comparisons; for example, states are merged when their merge score has exceeded the minimum score threshold. The algorithm ends when it cannot find any new merges. The final output is an approximation of the smallest state machine consistent with the input data.
Figure 2. An example of a PTA.
3. Assumptions of the Adversary
@@
Our work is grounded in a set of assumptions we formulated concerning the objectives and capabilities of an adversary. We consider an adversary attacking an arbitrary target network $TN$, driven by various objectives such as infiltrating $TN$ to find machines containing sensitive data, disrupting normal operations by launching denial of service attacks against specific services, orchestrating a botnet, or infecting machines with malware. In our main experiments, the training and test data from the datasets simulate an adversary attacking $TN$ without taking our defense into account. In our robustness experiments, we assume a more adept adversary who is aware of an ADS within $TN$ but unaware of its specifics. Their goal is to remain undetected until they have accomplished their objectives. To accomplish this, the adversary can monitor $TN$ to collect normal network behaviors and then attempt to bypass the ADS by generating behaviors resembling those collected during normal operations within $TN$. The adversary can only listen to the data that was sent but cannot delete or alter any of the sent data.
4. Related Works
@@
Traditional Methods. Several works have studied the efficacy of unsupervised learning for network anomaly detection. The primary focus has been on utilizing traditional unsupervised ML algorithms (e.g., Isolation Forest). Zoppi et al. conducted multiple studies evaluating the effectiveness of traditional unsupervised methods for anomaly detection on NetFlow data (Zoppi et al., 2021a, b). Campazas-Vega et al. also studied the effectiveness of Isolation Forests in detecting anomalies within NetFlow data. In contrast to SEQUENT, these studies do not dynamically adapt the anomaly scores based on the events/traces seen at test time for anomaly detection.
@@
State Machines. Numerous studies have investigated the efficacy of state machines for anomaly detection. Pellegrino et al. introduced BASTA, a system that learns probabilistic timed state machines from NetFlows and uses them as behavioral fingerprints for machines infected with botnet malware (Pellegrino et al., 2017). Alerts are raised when traces match behaviors captured in the fingerprints. Unlike SEQUENT, BASTA uses malicious data for learning rather than solely relying on benign data. Additionally, BASTA learns a state machine for each malware. This approach requires analysts to maintain an updated list of malware samples to learn fingerprints. Any missing malware samples may result in the system failing to detect new or unknown malware. SEQUENT learns a single state machine that models the behavior exhibited in a network and is not specific to any particular malware or network anomalies.
@@
Lin et al., Cao et al., and Matoušek et al. learned probabilistic state machines to detect anomalies in different systems(Cao et al., 2022; Lin et al., 2018; Matoušek et al., 2021). These works raise alerts when a new trace deviates significantly from what was observed during training. The degree of deviation is quantified based on the likelihood of a trace. In contrast, SEQUENT quantifies the abnormality based on state visit frequency for a corresponding trace and compares this to what was observed during training.
@@
Grov et al. presented a different approach for learning state machines to detect anomalies in NetFlows (Grov et al., 2019). The approach learns state machines using solely traces of network protocols extracted from benign data. The degree of abnormality is based on the normalized edit distance of an arbitrary trace under scrutiny; the number of changes needed such that each trace event can be replayed in the state machine. In contrast to SEQUENT, this work does not utilize the frequency of the states to adapt the scores dynamically at test time. Furthermore, it solely considers network protocols for the learning, whereas SEQUENT considers other features.
@@
Neural Networks. Deep Neural Networks (DNN) gained increasing popularity among researchers in recent years due to their ability to identify complex hidden patterns that traditional models may struggle to detect. Mirsky et al. introduced Kitsune, a state-of-the-art solution that utilizes an ensemble of Autoencoders (AE) to detect network anomalies in an unsupervised manner. It extracts various statistics from benign data and uses them for training. Kitsune then attempts to recreate instances it has seen during training by computing similar statistics for arbitrary new data. Alarms are raised for instances with large recreation errors(Mirsky et al., 2018). Neural networks are notoriously known to be difficult to interpret, as it is not always clear what happens between the input and output layers. Nguyen et al. sought to address this issue with their framework, GEE. It uses a Variational Autoencoder (VAE) to detect anomalies and generates explanations by providing SHAP-like plots of the gradients computed for NetFlow features(Nguyen et al., 2019). However, it is not clear how a security analyst should interpret these gradients. In contrast, SEQUENT learns an inherently interpretable model and generates root causes, which can be used to facilitate the understanding of (malicious) sequential behaviors. Root causes are linked back to concrete instances producing the anomalous behaviors.
5. SEQUENT
5.1. Intuition Behind SEQUENT
@@
SEQUENT employs a novel approach for computing the anomaly score of an arbitrary trace under scrutiny. Unlike the measures of abnormality presented in prior works (Pellegrino et al., 2017; Matoušek et al., 2021; Lin et al., 2018; Grov et al., 2019; Cao et al., 2022), SEQUENT leverages the frequencies of the state visits in a state machine to compute an adaptive anomaly score. The underlying principle of SEQUENT is that behaviors occurring significantly more often than those observed during training are indications of anomalies. Traces under scrutiny are replayed on a state machine to determine how well they fit within the learned model. Each trace visits a set of states, and the frequency of visits to each state is compared to the expected frequency observed during training. When the frequency of visits to a state is significantly higher than its expected frequency, it signals that unusual behavior is occurring in this state. The more significant the increase in visits for this state, the more likely it is to show anomalous behavior. SEQUENT employs this rationale to dynamically adapt its scores for arbitrary traces seen at test time. SEQUENT’s intuition will also work for (anomalous) behaviors never seen before during training, as these will visit unusual paths during the replay.
5.2. SEQUENT’s Anomaly Detection Pipeline
@@
Figure 3 depicts SEQUENT’s anomaly detection pipeline. It provides a high-level overview illustrating the flow of learning of a state machine from training data (represented by the blue arrows), the flow for running predictions on new data (represented by green arrows), and the flow for generating the results (represented by the orange arrows). Before NetFlows are used for training or prediction, they are first transformed into traces by the NetFlow Processor component, as depicted in the figure. Each flow is described in the following paragraphs.
Figure 3. SEQUENT’s anomaly detection pipeline.
@@
Processing NetFlows. NetFlow features need to be discretized before learning a state machine, given that each flow contains both continuous and discrete feature data. Furthermore, each flow needs to be transformed into an event (discrete symbol) to generate traces (sequence of events). Although the discretization process is not strictly necessary for the learning of a state machine, using using the raw NetFlow feature data (particularly continuous data) will lead to the combinatorial explosion of states. Discretization enables us to reduce the number of possible combinations in the feature data, thereby also reducing the number of states used in the state machine. To discretize NetFlow features, we develop an encoding algorithm inspired by the GloVe algorithm (Pennington et al., 2014). Our encoding algorithm computes co-occurrences between values for a given input flow feature $f$ and stores them in a matrix. Each row denotes a unique value $v$ of $f$ and how often it occurs with another value $v^{\prime}$. Feature values are discretized by first clustering the rows using KMeans and then using cluster assignments as their discretization label. We make our encoding algorithm available on Github (Cao, 2022a).
To transform a flow into an event symbol, we concatenate the discretized feature values of the selected features with an underscore symbol to form a single event symbol. For example, if three features are used to represent an event, we get the following event symbol: $DFV1\_DFV2\_DFV3$, where $DFV$ is the abbreviation for discretized feature value. In the case of one feature (e.g., bytes), we directly use its discretized values. To convert events into traces, we apply a sliding window of length $N$ to generate traces. These are then subsequently used to learn a state machine. Traces that do not have a length of $N$ are excluded from learning as they do not provide sufficient information.
@@
Learning of a State Machine. SEQUENT utilizes FlexFringe (Verwer and Hammerschmidt, 2017) to learn a state machine from NetFlow traces. This framework offers, out of the box, various heuristics for learning a state machine. Additionally, it enables its users to learn other types of models such as Markov Chains. FlexFringe also automates the creation of traces from the CSV file provided as input data. The user has to specify which features to use to create events, the length of the sliding window, and how the data should be sorted (e.g., by connections) to create the traces.
Since Flexfringe’s trace creation feature of FlexFrige matches the trace creation method mentioned in the previous paragraph, we leverage this feature of FlexFringe to generate traces from the provided input NetFlows. We opt for FlexFringe in the design of SEQUENT based on these aforementioned features.
Once the selected features of the training data are discretized, we forward the CSV containing discretized feature values of the training to FlexFringe to create traces and learn a state machine. We specify to FlexFringe to use the discretized features values for the creation of the event symbols.
When learning is completed, FlexFringe outputs a state machine along with various statistics, such as how often each state of the state machine was visited during training. We store the model and the statistics for the computation of the rolling anomaly score for traces observed at test time.
Computing Rolling Anomaly Scores. At test time, the new NetFlow data is processed in the similar manner as with the training data and is forwarded to FlexFringe. FlexFringe simulates each trace generated from the test data and computes the sequence of states visited by each trace. For each computed state sequence, we compute its rolling anomaly scores with the help of the statistics extracted during training of the model. Algorithm 1 depicts the pseudocode for computing rolling anomaly scores for an arbitrary trace $t$.
The first step in computing the rolling anomaly scores is to extract the state visit frequencies of each state and the sum of all visit frequencies (see lines 4 to 8). The sum of all visit frequencies (stored in “$TotalVisits$” from line 3) is used for normalization, as each state may have been visited a different number of times during training, causing some states to be more frequent than others. Note that lines 2 to 8 do not have to be repeated for each arbitrary trace we observe at test time. However, we have included it in Algorithm 1 for the sake of completeness. In our implementation of SEQUENT, we compute this once and store the result in memory.
The next step in computing rolling anomaly scores is to simulate the (test) trace $t$ on the state machine and extract the path it traverses in the model (line 10). For each state $q^{\prime}$ in the extracted state sequence (i.e., the path), we compute its anomaly score (lines 11 to 18). Note that since we are working with visit frequencies, large counts can have a considerable impact on the score. To address this, we use the logarithm of the frequencies to compute the score for $q^{\prime}$ (as shown on lines 12 to 17). The addition of one within the logarithm is used to ensure that we do not encounter the undefined behavior when of calculating the logarithm of zero. The base of the anomaly score for $q^{\prime}$ starts with its visit frequency observed so far at test time (line 12). If the $q^{\prime}$ has not been observed yet, a value of zero will be returned. The observed frequency (at test time) is then compared to the corresponding frequency extracted from the training statistics (line 13) and the score is normalized (lines 14 and 15). Finally, the visit frequency of state $q^{\prime}$ is updated (line 16), and the score is added to the results list (line 17). The anomaly score of trace $t$ is the sum of all individual scores in the result list. We keep the individual scores, as they are used to compute the root cause of an anomaly (explained in the next paragraph).
Input: State Machine $S$, Trace $t$, Training Statistics $TS$
Output: Rolling anomaly scores of $t$
1 $AnomalyScores\leftarrow[]$
2 $StateFreq_{train}\leftarrow\emptyset$
3 $TotalVisits\leftarrow 0$
4
5for  each state $q$ in $S$ do
6       $Freq_{q}\leftarrow extractVisitFreq(TS,q)$
7       $StateFreq_{train}[$$]\leftarrow freq_{q}$
8       $TotalVisit\leftarrow TotalVisit+Freq_{q}$
9
10 end for
11
[…]
13 $StateSequence\leftarrow computeStateSequence(S,t)$
14 for each state $q^{\prime}$ in $StateSequence$ do
15       $StateScore\leftarrow log(StateFreq_{test}[$q’]+1)]+1)] + 1 )
16       $StateScore\leftarrow StateScore-log(StateFreq_{train}[$q’]+1)]+1)] + 1 )
17       $StateScore\leftarrow StateScore+log(TotalVisit+1)$
18       $StateScore\leftarrow StateScore+log(|StateFreq_{test}|+1)$
[…]
20       $AnomalyScores.add(StateScore)$
21
22 end for
23
return AnomalyScores
Algorithm 1 SEQUENT’s Rolling Anomaly Score